OpenStudy (anonymous):

I need to find the derivative of (8x+9)^x. How do I approach that? Do I use the rule that g^x = g^x * ln (g) ?

OpenStudy (anonymous):

I mean that the derivative of g^x = g^x * ln *(g)

OpenStudy (turingtest):

yes, plus the chain rule...\[f(x)=(8x+9)^x\]chain rule:\[f'(x)=(8x+9)^x \ln (8x+9)(8)=8(8x+9)^x \ln (8x+9)\]

OpenStudy (anonymous):

The homework program in which I have to turn this in says this is false sadly. As a tip it gives taking the natural logarithm before deriving..

OpenStudy (turingtest):

Yeah, sorry. Seemed straight-forward to me. I'll try the log thing but in the meantime here's Wolfram's answer: http://www.wolframalpha.com/input/?i=d%2Fdx%28%288x%2B9%29^x%29

OpenStudy (turingtest):

\[y=(8x+9)^x\]\[\ln y=x \ln (8x+9)\]differentiating implicitly with the product rule gives\[y'/y=\ln(8x+9)+8x/(8x+9)\]\[y'=y[\ln(8x+9)+8x/(8x+9)]\]\[y'=(8x+9)^x[\ln(8x+9)+8x/(8x+9)]\]there we go!

OpenStudy (anonymous):

Thank you kindly good sir, just what I was looking for.

OpenStudy (turingtest):

anytime!